# QERA: an Analytical Framework for Quantization Error Reconstruction

### v1

@@ A-C1 QPEFT Hyperparameters
We sweep learning rates for each baseline/method, and collect the best accuracy. The reported accuracy is the average value across random seeds 42, 1, and 2. The total batch size is 64 for all GLUE experiments and we train the models for 5 epochs. For 4-bit experiments, we use 4-bit floating point from the QLoRA implementation in PEFT. For 3-bit experiments, we use emulated MXINT [37] with block size = 32 and for 2-bit experiments we use MXINT with block size = 16. Table V lists the learning rates for each experiment.
@@
We adopt the learning rates in meng2024pissa [38]. The reported perplexity/accuracy is the average value across random seeds 42, 1, and 2. For SlimPajama, we fine-tune the model on a subset for 1000 steps with rank = 8, total batch size = 64, sequence length = 1024, learning rate = 3e-5. For GSM8K, we fine-tune the model for 10 epochs with rank = 64, total batch size = 128, sequence length = 384, and learning rate = 3e-5.
TABLE V: Learning rates of RoBERTa-base experiments on GLUE.
@@ A-C2 PTQ Hyperparameters
A-D Decreasing Weight Error $\neq$ Decreasing Output Error for LoftQ
@@
A-E Test of Assumption 1
We provide more plots of normalized ${\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}}$ magnitude, $\frac{\mathrm{abs}(\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}})}{||\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}||_{F}}$, in Figures 7, 8, 9 and 10, where dark pixels are elements close to zeros. There are strongly correlated embedding channels in some k_proj and o_proj layers. The assumption fits better in MLP layers (gate_proj, up_proj, and down_proj).
A-F Choice of Calibration Set
@@
A-G Scalability and Numerical Stability of QERA
One may notice the diminishing model performance improvement of QERA-exact over QERA-approx as the model size increases. The main reason is that larger LLMs are more resistant to quantization [39]. Another reason can be the error ratio of the matrix square root calculation of the autocorrelation matrix increases with model hidden size (Figure 12(a)).
We find that the data type used in the calibration is important for the numeric stability of QERA-exact due to the calculation of the matrix square root and SVD. To improve the stability of the calculation in QERA-exact, a good practice we find is to perform the outer product of ${\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}$ in FP32, accumulated outer product in FP64, and calculate the matrix square root in FP64 using the blocked Schur algorithm [36]. Figure 12(b) illustrates the quantization time of QERA-approx and QERA-exact on the platform described in Section A-C where the linear layers are quantized sequentially. QERA-exact is slow due to the calculation of matrix square roots on CPUs. GPU-accelerated matrix square root will be the key optimization to reduce the quantization time. Note that in QERA, the quantization of individual layers is independent, allowing more parallelization and acceleration of the quantization process.
A-H Choices of LoRA Ranks and Models
@@
We notice LoftQ paper uses a large rank of 16 and 32 for fine-tuning on GLUE, which is larger than the commonly-used rank value of LoRA (4 or 8 in LoRA paper [12]). If we consider LoRA as the upper limit of QLoRA-like QPEFT methods (including LoftQ and QERA), to effectively compare these QPEFT methods, one easy way is to set the rank as the minimum value required by LoRA and check which QPEFT method achieves an accuracy closest to LoRA. This is why we choose rank = 8 for GLUE experiments (For 2-bit GLUE experiments we use a large rank 64 since the quantization is very aggressive). If we use rank = 32, LoRA and all the QPEFT methods may be over-parameterized and it will be hard to make a fair comparison in terms of fine-tuned accuracy.
@@
When investigating the related work, we find that both RoBERTa and DeBERTaV3 [40] are used in QPEFT experiments [3, 1, 38, 3, 41]. The reason why we chose RoBERTa is that the RoBERTa checkpoint on HuggingFace [^2] RoBERTa-base checkpoint: link is complete and compatible with both HuggingFace’s official examples of sequence classification [^3] HuggingFace example of sequence classification: link and masked language modeling [^4] HuggingFace example of masked language modeling: link. Specifically, the RoBERTa checkpoint contains both the base model and the masked language modeling head but the DeBERTaV3’s checkpoint [^5] DeBERTaV3’s checkpoint: link only contains the base model. As we know, the base model is enough for fine-tuning on downstream tasks. However, to calibrate on the pretraining dataset, we need the language modeling head to verify if our implementation of data preprocessing and calibration matches how the model was originally pretrained. Note that the quality of the statistic values in QERA like ${\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}}$ depends on the quality of the calibration set. Thus, without the language modeling head in the checkpoint, we cannot perform the QERA’s calibration for DeBERTaV3 properly, ensure the correctness of statistics in QERA, and explore the effect of the choice of calibration sets.
A-I Detailed PTQ Results
Here we offer the detailed evaluation results for each downstream task in Tables VI, VII, VIII, IX, X, XI and XII.
TABLE VI: Post-training quantization evaluation of TinyLlama-1.1B.
@@ A-I Detailed PTQ Results
TABLE VII: Post-training quantization evaluation of Gemma-2-2B.
@@
TABLE VIII: Post-training quantization evaluation of Phi3-3.5-mini.
@@
TABLE IX: Post-training quantization evaluation of LLaMA-2-7B.
@@
TABLE X: Post-training quantization evaluation of LLaMA-2-13B.
@@
TABLE XI: Post-training quantization evaluation of LLaMA-3.1-8B.
@@
TABLE XII: Post-training quantization evaluation of LLaMA-3.1-70B.

### v2

@@ A.4.1 QPEFT Hyperparameters
We sweep learning rates for each (method, task), and collect the best accuracy. Thus each (method, task) pair has its own tailored learning rate, ensuring the best performance of baselines and QERA under the same trainable parameter budget. The reported accuracy is the average value across random seeds 42, 1, and 2. The total batch size is 64 for all GLUE experiments and we train the models for 5 epochs. For 4-bit experiments, we use 4-bit floating point from the QLoRA implementation in PEFT. For 3-bit experiments, we use emulated MXINT (Darvish Rouhani et al., 2023) with block size = 32 and for 2-bit experiments we use MXINT with block size = 16. Table 6 lists the learning rates for each experiment.
@@
We adopt the learning rates in Meng et al. (2024). The reported perplexity/accuracy is the average value across random seeds 42, 1, and 2. For SlimPajama, we fine-tune the model on a subset for 1000 steps with rank = 8, total batch size = 64, sequence length = 1024, learning rate = 3e-5. For GSM8K, we fine-tune the model for 10 epochs with rank = 64, total batch size = 128, sequence length = 384, and learning rate = 3e-5.
Table 6: Learning rates of RoBERTa-base experiments on GLUE.
@@ A.4.2 PTQ Hyperparameters
A.5 Decreasing Weight Error $\neq$ Decreasing Output Error for LoftQ
@@
A.6 Choice of Calibration Set
@@
A.7 Scalability and Numerical Stability of QERA
One may notice the diminishing model performance improvement of QERA-exact over QERA-approx as the model size increases. The main reason is that larger LLMs are more resistant to quantization (Chee et al., 2024). Another reason can be the error ratio of the matrix square root calculation of the autocorrelation matrix increases with model hidden size (Figure 8(a)).
We find that the data type used in the calibration is important for the numeric stability of QERA-exact due to the calculation of the matrix square root and SVD. To improve the stability of the calculation in QERA-exact, a good practice we find is to perform the outer product of ${\bm{R}}_{{\mathbb{X}}{\mathbb{X}}}$ in FP32, accumulated outer product in FP64, and calculate the matrix square root in FP64 using the blocked Schur algorithm (Deadman et al., 2012). Figure 8(b) illustrates the quantization time of QERA-approx and QERA-exact on the platform described in Section A.4 where the linear layers are quantized sequentially. QERA-exact is slow due to the calculation of matrix square roots on CPUs. GPU-accelerated matrix square root will be the key optimization to reduce the quantization time. Note that in QERA, the quantization of individual layers is independent, allowing more parallelization and acceleration of the quantization process.
A.8 Choice of Solutions for QPEFT and PTQ
QPEFT and PTQ are two different application scenarios of QERA. We recommend QERA-approx for QPEFT and QERA-exact for PTQ. PTQ aims to recover the model performance as much as possible without re-training. For PTQ, it is desirable to recover more model performance even if it takes longer to compute low-rank terms. Note that the low-rank terms are pre-computed once offline. At inference time, QERA-exact introduces no overhead to the hardware since LQER, QERA-approx, and QERA-exact all takes the same form of ${\bm{y}}={\bm{x}}(\widetilde{{\bm{W}}}+{\bm{A}}_{k}{\bm{B}}_{k})$.
However, for QPEFT experiments, it is unreasonable to pay a long time for initializing the low-rank terms for the limited improvement in output approximation error (i.e., QERA-exact/CALDERA), because 1) fine-tuning can recover the error, and 2) instead of spending much time on initialization, increasing training steps or increasing the rank number brings more gain in the fine-tuned accuracy. We run controlled experiments to support this claim. In Table 7 and Table 8, we run QPEFT experiments of RoBERTA-base on MRPC and LLaMA-2-7B on SlimPajama respectively. Compared to QERA-exact (Caldera’s Lemma 4.2), QERA-approx achieves better accuracy/perplexity while taking $\frac{2}{3}\sim\frac{1}{2}$ of the time.
Table 7: Runtime comparison of QERA-exact and QERA-approx on MRPC. It is recommended using QERA-approx for QPEFT instead of QERA-exact.
Method Rank Epochs Init. time Training time Total time ($\downarrow$) Acc. ($\uparrow$)
QERA-exact 8 4 1.6min 2.2min 3.8min 88.97
QERA-approx 12 4 21s 2.2min 2.6min 89.95
QERA-approx 8 5 21s 2.7min 3.1min 89.97
Table 8: Runtime comparison of QERA-exact and QERA-approx on SlimPajama. It is recommended using QERA-approx for QPEFT instead of QERA-exact.
Method Rank Epochs Init. time Training time Total time ($\downarrow$) PPL. ($\downarrow$)
QERA-exact 16 2 4.9h 1.9h 6.8h 6.31
QERA-approx 64 2 29.6min 2.1h 2.6h 6.18
QERA-approx 16 4 28.2min 4.0h 4.5h 6.21
A.9 Choices of LoRA Ranks, Models, and Precisions for QPEFT
@@
We notice LoftQ paper uses a large rank of 16 and 32 for fine-tuning on GLUE, which is larger than the commonly-used rank value of LoRA (4 or 8 in LoRA paper (Hu et al., 2021)). If we consider LoRA as the upper limit of QLoRA-like QPEFT methods (including LoftQ and QERA), to effectively compare these QPEFT methods, one easy way is to set the rank as the minimum value required by LoRA and check which QPEFT method achieves an accuracy closest to LoRA. This is why we choose rank = 8 for GLUE experiments (For 2-bit GLUE experiments we use a large rank 64 since the quantization is very aggressive). If we use rank = 32, LoRA and all the QPEFT methods may be over-parameterized and it will be hard to make a fair comparison in terms of fine-tuned accuracy. To support this claim, we sweep the rank of LoRA-adapted RoBERTA-base on SST2 and MRPC and show a large rank $k$ like 16 in LoftQ has over-parallelization problem in Table 9 and Table 10.
@@
When investigating the related work, we find that both RoBERTa and DeBERTaV3 (He et al., 2021) are used in QPEFT experiments (Guo et al., 2023; Li et al., 2023; Meng et al., 2024; Guo et al., 2023; Zhang et al., 2023). The reason why we chose RoBERTa is that the RoBERTa checkpoint on HuggingFace [^2] RoBERTa-base checkpoint: link is complete and compatible with both HuggingFace’s official examples of sequence classification [^3] HuggingFace example of sequence classification: link and masked language modeling [^4] HuggingFace example of masked language modeling: link. Specifically, the RoBERTa checkpoint contains both the base model and the masked language modeling head but the DeBERTaV3’s checkpoint [^5] DeBERTaV3’s checkpoint: link only contains the base model. As we know, the base model is enough for fine-tuning on downstream tasks. However, to calibrate on the pretraining dataset, we need the language modeling head to verify if our implementation of data preprocessing and calibration matches how the model was originally pretrained. Note that the quality of the statistic values in QERA like ${\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}}$ depends on the quality of the calibration set. Thus, without the language modeling head in the checkpoint, we cannot perform the QERA’s calibration for DeBERTaV3 properly, ensure the correctness of statistics in QERA, and explore the effect of the choice of calibration sets.
Table 9: Over-parameterization problem. We sweep the rank $k$ of LoRA on SST2 and reported fine-tuned accuracy. The highest accuracy at rank $k=12$ indicates over-parameterization happens for $k\geq 12$.
Method Rank $k$ Learning rates Best Acc.
LoRA 4 1e-4/2e-4/3e-4/4e-4/5e-4/6e-4 94.38
8 1e-4/2e-4/3e-4/4e-4/5e-4/6e-4 94.46
12 1e-4/2e-4/3e-4/4e-4/5e-4/6e-4 94.73
16 1e-4/2e-4/3e-4/4e-4/5e-4/6e-4 94.50
20 1e-4/2e-4/3e-4/4e-4/5e-4/6e-4 94.50
Table 10: Over-parameterization problem. We sweep the rank $k$ of LoRA on MRPC and reported fine-tuned accuracy. The highest accuracy at rank $k=12$ indicates over-parameterization happens for $k\geq 12$.
Method Rank $k$ Learning rates Best Acc.
LoRA 4 1e-4/2e-4/3e-4/4e-4/5e-4/6e-4 87.99
8 1e-4/2e-4/3e-4/4e-4/5e-4/6e-4 88.97
12 1e-4/2e-4/3e-4/4e-4/5e-4/6e-4 89.95
16 1e-4/2e-4/3e-4/4e-4/5e-4/6e-4 89.46
20 1e-4/2e-4/3e-4/4e-4/5e-4/6e-4 89.71
A.10 Detailed PTQ Results
Here we offer the detailed evaluation results for each downstream task in Tables 11, 12, 13, 14, 15, 16 and 17.
Table 11: Post-training quantization evaluation of TinyLlama-1.1B.
@@ A.10 Detailed PTQ Results
Table 12: Post-training quantization evaluation of Gemma-2-2B.
@@
Table 13: Post-training quantization evaluation of Phi3-3.5-mini.
@@
Table 14: Post-training quantization evaluation of LLaMA-2-7B.
@@
Table 15: Post-training quantization evaluation of LLaMA-2-13B.
@@
Table 16: Post-training quantization evaluation of LLaMA-3.1-8B.
@@
Table 17: Post-training quantization evaluation of LLaMA-3.1-70B.
@@
A.11 Test of Assumption 1
We provide more plots of normalized ${\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}}$ magnitude, $\frac{\mathrm{abs}(\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}})}{||\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}||_{F}}$, across LLaMA-3.1-8B, LLaMA-2-7B, Mistral-7B-v0.3, and TinyLlama-1.1B in Figures 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23 and 24, where dark pixels are elements close to zeros. There are strongly correlated embedding channels in some k_proj and o_proj layers. The assumption fits better in MLP layers (gate_proj, up_proj, and down_proj), and holds for over 60% of the layers in LLMs.
Figure 13: Normalized $\mathrm{abs}({\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}})$ of inputs of k_proj layers in LLaMA-2-7B. Note that the q_proj and v_proj share the same inputs. Layers are sampled and only the first 96 dimensions are plotted for clarity.
Figure 14: Normalized $\mathrm{abs}({\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}})$ of inputs of o_proj layers in LLaMA-2-7B. Layers are sampled and only the first 96 dimensions are plotted for clarity.
Figure 15: Normalized $\mathrm{abs}({\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}})$ of inputs of gate_proj layers in LLaMA-2-7B. Note that the up_proj shares the same inputs. Layers are sampled and only the first 96 dimensions are plotted for clarity.
Figure 16: Normalized $\mathrm{abs}({\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}})$ of inputs of down_proj layers in LLaMA-2-7B. Layers are sampled and only the first 96 dimensions are plotted for clarity.
Figure 17: Normalized $\mathrm{abs}({\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}})$ of inputs of k_proj layers in Mistral-7B-v0.3. Note that the q_proj and v_proj share the same inputs. Layers are sampled and only the first 96 dimensions are plotted for clarity.
Figure 18: Normalized $\mathrm{abs}({\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}})$ of inputs of o_proj layers in Mistral-7B-v0.3. Layers are sampled and only the first 96 dimensions are plotted for clarity.
Figure 19: Normalized $\mathrm{abs}({\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}})$ of inputs of gate_proj layers in Mistral-7B-v0.3. Note that the up_proj shares the same inputs. Layers are sampled and only the first 96 dimensions are plotted for clarity.
Figure 20: Normalized $\mathrm{abs}({\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}})$ of inputs of down_proj layers in Mistral-7B-v0.3. Layers are sampled and only the first 96 dimensions are plotted for clarity.
Figure 21: Normalized $\mathrm{abs}({\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}})$ of inputs of k_proj layers in TinyLlama-1.1B. Note that the q_proj and v_proj share the same inputs. Layers are sampled and only the first 96 dimensions are plotted for clarity.
Figure 22: Normalized $\mathrm{abs}({\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}})$ of inputs of o_proj layers in TinyLlama-1.1B. Layers are sampled and only the first 96 dimensions are plotted for clarity.
Figure 23: Normalized $\mathrm{abs}({\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}})$ of inputs of gate_proj layers in TinyLlama-1.1B. Note that the up_proj shares the same inputs. Layers are sampled and only the first 96 dimensions are plotted for clarity.
Figure 24: Normalized $\mathrm{abs}({\mathbb{R}_{{\mathbb{X}}{\mathbb{X}}}})$ of inputs of down_proj layers in TinyLlama-1.1B. Layers are sampled and only the first 96 dimensions are plotted for clarity.